# RGS-DR: Reflective Gaussian Surfels with Deferred Rendering for Shiny Objects

### v3

@@ 5.2 Implementation Details
All our models are trained on an NVIDIA Tesla P100 GPU (16GB VRAM), completing 35k iterations in roughly two hours. For the lighting pass, we use a cube mipmap of size $6\times 512\times 512$ with 3 RGB channels, and for the residual pass an $8\times 512\times 512$ mipmap with 16 feature channels, along with a shallow MLP (two hidden layers of 256 units). For the real scenes, we downsample the training images of the gardenspheres and toycar scenes by factors of 4 and the sedan scene by a factor of 8 to be consistent with 3DGS-DR [16]. Furthermore, we employ a similar bounding volume as in [16] to allow for reflective surfaces only in the foreground object and penalize reflective surfels in the background. We achieve this by penalizing low roughness properties in Gaussians outside of the bounding volume. Loss weights are set as follows: normal consistency loss ($\lambda_{n}=0.05$) begins at 5k iterations, depth distortion loss ($\lambda_{d}=100$) at 3k iterations, and alpha loss ($\lambda_{\alpha}=1.0$) from the start—downweighted by a factor of 0.8 at 7k iterations. Material properties (diffuse color, roughness, tint) are optimized at a learning rate of $2.5\times 10^{-3}$, while opacity, scaling, and rotation use rates of $0.03$, $3\times 10^{-3}$, and $10^{-2}$, respectively. The environment cube mipmap follows an exponential decay from $10^{-2}$ to $10^{-3}$, and the positions of the 2D Gaussian discs are trained following the original 2DGS methodology [11] with a decay from $1.6\times 10^{-4}$ to $1.6\times 10^{-6}$. Initially, the model is trained without the residual pass. After 30k iterations, we activate it while freezing the geometry, material properties, and lighting model. At this stage, only the spherical encoding mipmap, MLP weights, and features are updated, so the residual term benefits mainly reconstruction and not relighting.
@@ 5.3 Comparison
Table 1: Quantitative results comparing our method with state-of-the-art approaches on test views from the Shiny Synthetic and Glossy Synthetic datasets. Metrics from [39, 16] are used, with the top three methods for each scene and metric marked as first, second, and third. Methods indicated by $\dagger$ do not support relighting. The final columns present the average scores for each dataset.
Shiny Synthetic [33] Glossy Synthetic [19] Shiny Real [33]
Datasets ball car coffee helmet teapot toaster avg bell cat luyu potion tbell teapot avg garden sedan toycar avg
@@
In Fig. 5, we present the decomposition of our method’s output including the material properties, smooth geometry (visualized through surface normals), as well as the estimated specular components and residuals. As demonstrated, our method effectively disentangles the geometry, material properties, and illumination of a given scene, thus enabling photorealistic novel view synthesis, relighting, and scene editing. Unlike 3DGS-DR, our method learns explicit material properties and thus trained scenes can be easily ported to existing 3D rendering engines.
@@
In Fig. 6, we present our method’s estimated environment illumination for four scenes of the Shiny Synthetic dataset. Our method manages to effectively disentangle the illumination from the material properties and geometry of the scene. Compared to 3DGS-DR and GaussianShader, we obtain more refined environment maps with sharper details, fewer artifacts, and closer to the ground truth.
Figure 6: Estimated environment maps for four synthetic scenes from our method RGS-DR, 3DGS-DR [16] and GaussianShader [12].
@@ 5.4 Ablation Study
In Table 2, we conduct an ablation study to showcase the contributions of our method. First, we replace the cube mipmap of the environment light with a single cubemap to compare with 3DGS-DR that uses a single level. We observe a considerable drop in performance, demonstrating the necessity of a multi-resolution cubemap for accurate estimates of environment illumination. Second, we ablate the residual pass and once again notice a drop in performance, which highlights its benefit to photorealistic reconstruction.
@@ 5.5 Relighting and Scene Editing
In Fig. 7, we present the relighting capabilities of our model for three scenes from the Shiny Synthetic dataset [33] and three scenes from the Glossy Synthetic dataset [19]. To relight the scenes, we apply gamma tonemapping to the new environment maps, replace the learned environment map, and convert them to a cube mipmap before rendering. As demonstrated, our method can effectively relight learned scenes and achieve photorealistic results even under new illuminations.
Figure 7: Relighted scenes. The first column displays the original rendered images while the subsequent columns show relighted scenes with their corresponding environment maps above.
In Fig. 8, on the other hand, we demonstrate the scene editing capabilities of our model on three glossy objects. Specifically, we apply roughness and diffuse editing to make the objects more or less reflective or change their diffuse color.
Figure 8: Scene editing: from left to right, the original rendering is followed by versions with decreased roughness, increased roughness, and a modified diffuse color.
@@ 6 Conclusion
This work addresses the challenges of reconstructing and performing inverse rendering on glossy objects exhibiting intricate geometries and complex view-dependent effects. Our RGS-DR leverages 2D Gaussian Splatting to efficiently achieve precise geometric reconstruction while simultaneously optimizing material properties and scene illumination. By introducing a new deferred shading paradigm that merges a conventional deferred lighting pass with a residual pass to fill in missing details, we significantly enhance the reconstruction quality, achieving performance on par with—and often surpassing—methods dedicated exclusively to reconstruction. Extensive experiments on challenging datasets featuring highly reflective objects demonstrate that our method not only competes with state-of-the-art techniques in reconstruction and inverse rendering, but also delivers high-quality relighting and scene editing.
@@ Acknowledgments
The resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation - Flanders (FWO) and the Flemish Government.
@@
In this supplementary material, we provide additional qualitative results to showcase the contributions of our proposed methodology, including model outputs and scene decompositions, estimated environment maps, relightings. Last but not least, we provide a video demonstration of our method with smooth animations of rendered scenes showcasing the capabilities of our method for reconstruction, relighting and scene editing. We compare our results with the corresponding results of 3DGS-DR [16] and Gaussian Shader [12] and demonstrate the superiority of our approach.
Appendix A Model Outputs and Scene Properties
@@
Appendix B Recovered Environment Maps
In Fig. 11 and Fig 12 we show additional results on estimated environment maps for all evaluated scenes on the Shiny Synthetic [33] and Glossy Synthetic [19] datasets. As evident, our method consistently outperforms 3DGS-DR [16] and GaussianShader [12], on the shinier scenes e.g. car, helmet and the entire Glossy Synthetic scenes, while being competitive on the rest. In general, we recover sharper details of the environment illumination with less artifacts or blurring, although we note a drop in quality for the rougher objects, e.g. coffee and teapot from the Shiny Synthetic dataset.
@@ Appendix B Recovered Environment Maps
Appendix C Additional Relighting Results
In Fig. 13 and Fig. 14, we demonstrate the relighting capabilities of our method by applying three novel environment illuminations on all examined scenes from the two synthetic datasets. We observe high-quality relighted reconstructions with preserved underlying geometry and material properties, while under different lighting.
@@ Appendix C Additional Relighting Results
Figure 14: Relighting results on the entire set of examined scenes from the Glossy Synthetic dataset [19]. High-quality and sharp reflections and specular highlights are observed in all scenes and especially the more glossy scenes, showcasing the successful estimate of scene properties. From left to right, we present the optimized scenes with their original environment illumination, followed by three relighting experiments per scene using three distinct environment maps, namely bridge, courtyard, and fireplace.
Appendix D Supplementary Video
In the supplementary video, we demonstrate the outputs of our method using smooth animations of reconstructed scenes. First, we showcase the decomposition of a scene into its material properties and light components (diffuse, specular, residual), as well as the underlying geometry (surface normals). Then we show side-by-side comparisons of rendered objects by our method versus 3DGS-DR [16] and GaussianShader [12]. Our method clearly outperforms both methods on geometry estimation thanks to the utilized 2D Gaussian surfels, showcasing less artifacts, smoother geometry, better appearance reconstruction than GaussianShader and usually on par or better reconstruction than 3DGS-DR. Then we present some additional relighting results before ending the video with more high-quality rendered animations of glossy objects. The video can be watched at the following URL: https://drive.google.com/file/d/1gyIV5-40NtyH9IKnQSRiwRiur7RFF3pM/view?usp=sharing
Appendix E Limitations
Our results reveal several limitations. First, the assumption of deferred shading that each pixel corresponds to the nearest opaque surface prevents proper handling of transparent or refractive objects — a shortcoming that might be overcome by combining deferred rendering for opaque surfaces with forward rendering for transparent ones. Second, the absence of multi-bounce ray tracing and explicit visibility reasoning hinders our ability to manage interreflections and shadows, adversely affecting performance on the coffee and teapot scenes of the Shiny Synthetic dataset [33]. Third, the residual pass is only used for refining the model for novel view synthesis without any benefit to the geometry estimation or relighting tasks. Ideally, the residual should be used to compensate for interreflections and other poorly modelled effects during optimization to alleviate geometry distortions as well. Finally, the approach is not yet applicable to unbounded scenes, requiring either ray contraction [1] and a bounding volume around the foreground object [16] to mitigate interference from sparsely observed background features. Such sparsely seen features in the background are perceived as view-dependent effects that result in erroneous reflective surfels. At the moment, this can only be addressed by using a bounding volume that penalizes low roughness of all surfels outside the bounding volume.

### v4

@@ 5.2 Implementation Details
All our models are trained on an NVIDIA Tesla P100 GPU, completing 35k iterations in two hours on average. For the lighting pass, we use a cube mipmap of size $6\times 512\times 512$ with 3 RGB channels, and for the residual pass an $8\times 512\times 512$ mipmap with 16 feature channels, along with a shallow MLP (two hidden layers of 256 units) plus one 4D feature per Gaussian. For the real scenes, we downsample the training images of the gardenspheres and toycar scenes by factors of 4 and the sedan scene by a factor of 8 to be consistent with [16, 43]. Furthermore, we employ a similar bounding volume as in [16, 43] to allow for reflective surfaces only in the foreground object. We achieve this by penalizing low roughness properties in Gaussians outside of the bounding volume. We use loss weights of $\lambda_{n}=0.05$, $\lambda_{d}=100$, and $\lambda_{\alpha}=1.0$. Material properties are optimized with a learning rate of $2.5\times 10^{-3}$, while opacity, scale, and rotation use $0.03$, $3\times 10^{-3}$, and $10^{-2}$, respectively. The environment cube mipmap is updated with an exponentially decaying rate from $10^{-2}$ to $10^{-3}$. Gaussian positions are trained as in 2DGS [11], with a decay from $1.6\times 10^{-4}$ to $1.6\times 10^{-6}$. We train 30k iters without the residual, then enable it and freeze geometry/materials/lighting for a 5k-iter fine-tune. Only the residual branch is updated, benefiting NVS but not affecting editing.
@@ 5.3 Comparison
Table 1: Quantitative results comparing our method with state-of-the-art approaches on test views from the Shiny Synthetic, Shiny Real, and Glossy Synthetic datasets. Metrics from [43, 16] are used, with the top three methods for each scene and metric marked as first, second, and third. The final columns present the average scores for each dataset. The second column indicates whether a method supports relighting (✓) or is only reconstruction-capable (✗) .
Rel- Shiny Synthetic [35] Glossy Synthetic [20] Shiny Real [35]
ight ball car coffee helmet teapot toaster avg bell cat luyu potion tbell teapot avg garden sedan toycar avg
@@
In Fig. 4, we present the decomposition of our method’s output, including the material properties, smooth geometry (visualized through surface normals), as well as the estimated specular components and residuals. As demonstrated, our method effectively disentangles the geometry, material properties, and illumination of a given scene, thus enabling photorealistic novel view synthesis, relighting, and scene editing. Unlike 3DGS-DR, our method learns explicit material properties and thus trained scenes can be easily ported to existing 3D rendering engines.
@@
In Fig. 5 and Fig. LABEL:fig:real-envmaps, we present our method’s estimated environment illumination for scenes from the Shiny Synthetic and Shiny Real datasets. Our method manages to effectively disentangle the illumination from the material properties and geometry of the scene. Compared to 3DGS-DR and GaussianShader, we obtain more refined environment maps with sharper details and fewer artifacts. A quantitative evaluation is provided in Table 2, containing the reconstruction error (MAE) and environment map quality (E-PSNR, E-SSIM, E-LPIPS) for the Shiny Synthetic dataset.
Table 2: Quantitative comparison of geometry reconstruction and environment illumination on the Shiny Blender dataset.
MAE $\downarrow$ E-PSNR $\uparrow$ E-SSIM $\uparrow$ E-LPIPS $\downarrow$
GShader 8.76 11.06 0.142 0.644
3DGS-DR 4.95 11.89 0.339 0.599
Ours 4.24 13.28 0.403 0.563
Figure 5: Comparison of estimated environment maps of synthetic scenes from the Shiny Synthetic dataset [35].
Figure 6: Relighted scenes with GT on the left and utilized environment map on the top.
@@ 5.4 Ablation Study
In Table 3, we conduct an ablation study to validate our design choices. First, we replace the cube mipmap of the environment light with a single cubemap to compare with 3DGS-DR which uses a single level. We observe a considerable drop in performance, demonstrating the necessity of a multi-resolution cubemap for accurate estimates of environment illumination. Second, we ablate the residual pass and once again notice a drop in performance, which highlights its benefit to photorealistic NVS. Finally, we replace the residual pass with spherical-harmonics-based residuals like in [12], which improves performance but not as much as our proposed residual pass.
@@ 5.5 Relighting and Scene Editing
In Fig. 6, we present the relighting capabilities of our model for two scenes from the Shiny Synthetic dataset [35] and two scenes from the Glossy Synthetic dataset [20]. To relight the scenes, we replace the learned environment map and convert it to a cube mipmap before rendering. As demonstrated, our method can effectively relight learned scenes and achieve photorealistic results even under new illuminations. Additional relighting results can be found in the supplementary material. In Fig. 7, on the other hand, we demonstrate the scene editing capabilities of our model on three glossy objects by modifying the roughness to make them less or more reflective and changing the diffuse color.
Figure 7: Scene editing: original rendering, decreased roughness, increased roughness, and modified diffuse color.
@@ 6 Conclusion
This work tackles inverse rendering of glossy objects with complex view-dependent effects using a 2D Gaussian, pixel-deferred formulation. Our method estimates editable material properties and environment illumination via split-sum IBL, and adds a refinement stage with a lightweight directional residual that captures leftover specular detail, further enhancing novel view synthesis and bridging the gap with reconstruction-only methods. Across challenging reflective scenes, we observe sharper highlights, plausible material estimates, and higher relighting fidelity than per-Gaussian shading baselines or uniform roughness baselines. Our scope is limited to direct IBL since we do not model multi-bounce indirect lighting, transparency, or subsurface scattering. Scenes dominated by strong near-field interreflections are better addressed by global-illumination-enabled variants.
@@ Acknowledgments
This work is supported by the Flanders AI Research Program. The resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation - Flanders (FWO) and the Flemish Government.
@@
In this supplementary material, we provide additional qualitative results to showcase the contributions of our proposed methodology, including model outputs and scene decompositions, estimated environment maps, and relightings. Last but not least, we provide a video demonstration of our method with smooth animations of rendered scenes showcasing the capabilities of our method for reconstruction, relighting and scene editing. We compare our results with the corresponding results of 3DGS-DR [16] and Gaussian Shader [12] and demonstrate the superiority of our approach.
7 Model Outputs and Scene Properties
@@
8 Recovered Environment Maps
In Fig. 10, Fig 11, and Fig. 12, we show additional results on estimated environment maps for all evaluated scenes on the Shiny Synthetic [35] and Glossy Synthetic [20] datasets. As evident, our method consistently outperforms 3DGS-DR [16] and GaussianShader [12], on the shinier scenes e.g. car, helmet, and the entire Glossy Synthetic scenes, while being competitive on the rest. In general, we recover sharper details of the environment illumination with fewer artifacts or blurring.
@@ 8 Recovered Environment Maps
Figure 12: Estimated environment maps on the Shiny Real dataset [20] recovered by our method RGS-DR, 3DGS-DR [16] and GaussianShader [12]. In all cases, we recover sharper details with less blurriness or noise than the competing methods.
9 Additional Relighting Results
In Table 4 and Fig. 13, we present a comparison of our method against GaussianShader [12] and 3DGS-DR [16] on relighting of the Glossy Synthetic [20] scenes, demonstrating the superiority of our method both quantitatively and qualitatively. We provide additional qualitative relighting results of our method in Fig. 14 for the Shiny Synthetic scenes [35] and Fig. 15 for the Glossy Synthetic scenes [20] under three distinct environment maps.
Table 4: Comparison of relighting performance on the Glossy Synthetic [20] dataset.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
3DGS-DR 20.541 0.887 0.091
GShader 24.236 0.906 0.089
Ours 24.474 0.920 0.070
Figure 13: Comparison of relighting results on the Glossy Synthetic scenes [20] between our method, GaussianShader [12] and 3DGS-DR [16]. As can be easily observed, our relighting results are closer to the GT, with sharper reflections and specular highlights.
@@ 9 Additional Relighting Results
Figure 15: Relighting results on the entire set of examined scenes from the Glossy Synthetic dataset [20]. High-quality and sharp reflections and specular highlights are observed in all scenes and especially in the more glossy scenes, showcasing the successful estimate of scene properties. From left to right, we present the optimized scenes with their original environment illumination, followed by three relighting experiments per scene using three distinct environment maps, namely bridge, courtyard, and fireplace.
10 Supplementary Video
In the supplementary video  [^1] https://drive.google.com/file/d/1tJBisVljfYMnsNNUS7qHZ4TerOcyqJfM/view?usp=sharing, we demonstrate the outputs of our method using smooth animations of reconstructed scenes. First, we showcase the decomposition of a scene into its material properties and light components (diffuse, specular, residual), as well as the underlying geometry (surface normals). Then we show side-by-side comparisons of rendered objects by our method versus 3DGS-DR [16] and GaussianShader [12]. Our method clearly outperforms both methods on geometry estimation thanks to the utilized 2D Gaussian surfels, showcasing less artifacts, smoother geometry, better appearance reconstruction than GaussianShader and usually on par or better reconstruction than 3DGS-DR. We then present some additional relighting results before concluding the video with more high-quality, rendered animations of glossy objects.
11 Limitations
Our results reveal several limitations. First, the deferred shading assumption, that each pixel corresponds to the nearest opaque surface, prevents accurate handling of transparent or refractive materials. This could potentially be addressed by combining deferred rendering for opaque regions with forward rendering for transmissive ones. Second, the lack of multi-bounce ray tracing and explicit visibility reasoning limits our ability to capture interreflections and shadows, leading to degraded performance on scenes such as coffee and teapot from the Shiny Synthetic dataset [35]. Recent methods [7, 39, 38] incorporate indirect light transport into direct IBL, making them better suited for such complex lighting conditions. Third, our residual pass is only used to improve novel view synthesis and does not contribute to geometry estimation or relighting. Ideally, it could also be leveraged during optimization to compensate for interreflections and other hard-to-model effects, potentially reducing geometric artifacts. Nonetheless, its current role is primarily to narrow the gap with reconstruction-focused methods without explicitly modeling higher-order light transport.